on complete combustion 0.246g of an organic compound gave 0.198g of CO2 and 0.1014g of H2O. What is the ratio of carbon and hydrogen atoms in the compound?


1
Expert's answer
2022-03-15T11:42:05-0400

Solution:

The molar mass of CO2 is 44.01 g/mol

Therefore,

Moles of CO2 = (0.198 g CO2) × (1 mol CO2 / 44.01 g CO2) = 0.00450 mol CO2


The molar mass of H2O is 18.0153 g/mol

Therefore,

Moles of H2O = (0.1014 g H2O) × (1 mol H2O / 18.0153 g H2O) = 0.00563 mol H2O


Schematic equation:

CO2 → C

According to the shematic equation above:

Moles of C = Moles of CO2 = 0.00450 mol

Moles of C = 0.00450 mol


Schematic equation:

H2O → 2H

According to the shematic equation above

Moles of H = 2 × Moles of H2O

Moles of H = 2 × 0.00563 mol = 0.01126 mol

Moles of H = 0.01126 mol


Thus, the ratio of carbon and hydrogen atoms in the compound is:

Moles of C : Moles of H = 0.00450 mol : 0.01126 mol = 1 : 2.5

Moles of C : Moles of H = 1 : 2.5


Answer: The ratio between the carbon and hydrogen atoms is 1 : 2.5